The derivative of y = ln 3x 4


1
Expert's answer
2021-07-05T17:47:01-0400

y=ln(3x+4)

dy/dx=d/dx(ln(3x+4))

dy/dx=1/(3x+4) * d/dx(3x)

=3/(3x+4)

Answer